# TTA-Nav: Test-time Adaptive Reconstruction for Point-Goal Navigation under Visual Corruptions

### 04 Mar 2024 v1

@@ Abstract
Robot navigation under visual corruptions presents a formidable challenge. To address this, we propose a Test-time Adaptation (TTA) method, named as TTA-Nav, for point-goal navigation under visual corruptions. Our “plug-and-play” method incorporates a top-down decoder to a pre-trained navigation model. Firstly, the pre-trained navigation model receives a corrupted image input and extracts features. Secondly, the top-down decoder produces the reconstruction given the high-level features extracted by the pre-trained model. Then, it feeds the reconstruction of a corrupted image back to the pre-trained model, which outputs the final action. Despite being trained solely on clean images, the top-down decoder reconstructs cleaner images from corrupted ones without gradient-based adaptation. The pre-trained navigation model with our top-down decoder significantly enhances navigation performance across almost all visual corruptions in our benchmarks. Our method improves the success rate of point-goal navigation from the state-of-the-art result of 46% to 94% on the most severe corruption, showing the potential for broader application in robotic visual navigation.
@@ I Introduction
Robot navigation faces substantial challenges when confronted with visual corruptions [1, 2, 3, 4, 5], as illustrated in Fig. 1: The robot is equipped with a state-of-the-art (SOTA) navigation model trained on 2.5 billion clean frames with deep reinforcement learning (DRL). Despite being trained on a great amount of data, the robot still fails to navigate to the target position under dimmed lighting.
@@
This challenge raises questions about the adaptability of robots: How can a robot dynamically adjust in real-time to visual corruptions that are unforeseeable in the training phase? And how can it maintain effective navigation when encountering such corruption despite being trained solely on clean observations? The previous work [5] considers the navigation with a map, which can be deteriorated if the camera images are corrupted, and proposed a map refiner to generate a more precise map. However, their method requires the robot to collect data and gradient-based fine-tuning in the test time. That means the robot has a distinct fine-tuning and test phase. This is impractical in some applications that require the robot to immediately adapt to corruption, such as self-driving cars.
@@
Figure 1: Point-Goal Navigation Under Visual Corruptions: A robot is tasked to move from start to goal position. Top: The robot is equipped with the state-of-the-art navigation method [6]. It fails to navigate from the bedroom to another room when facing the dimmed light condition. Bottom: The robot is equipped with our method. It gets the reconstructed image as the “surrogate observation” instead of the real observation from the scene.
@@
Figure 2: An overview of our proposed method (TTA-Nav): We present the pretrained navigation model (Pretrained-Nav), which contains Visual Encoder (VE) and policy network, in the purple frame. TTA-Nav is a plug-and-play method. It has Top-down Decoder (TD) in the orange frame. TD receives the output of the late layer of VE and projects the reconstructed image back to the VE’s input layer. Here VE is SE-ResNeXt-50 [7] and TD receives block3’s output. Left: During training, VE remains fixed, while TD is trained to predict VE’s input by minimizing Mean Squared Error (MSE) loss. Training occurs offline, utilizing samples from Replay, which contains the visual experiences of a robot navigating 72 Gibson training scenes. Right: During testing, TD is frozen and the BatchNorm layers of VE can adjust their normalization statistics $\hat{\mu}_{k}$ and $\hat{\sigma}_{k}$. See section III for more details. Finally, the reconstructed image is intput to VE for navigation. VE does feedforward computation again and the policy outputs actions.
@@
To this end, we propose a Test-time Adaptation (TTA) method for robot navigation under visual corruptions. Our method is inspired by the top-down processing observed in human cognition [8]. This processing is deemed as a fundamental mechanism for solving challenging visual tasks [9, 10]. Specifically, consider a Deep Neural Network (DNN) model as an architecture with low-level visual processing at earlier layers and high-level visual processing at later layers, our method relies on Top-down modulation from a Decoder, which receives a latent low-level representation of the camera review from the pre-trained visual perception model, reconstructs the clean observation and feeds reconstructed inputs back to the navigation model. The other key mechanism of our method is adaptive normalization (AN) in the pre-trained visual perception model. During the test phase, AN adjusts normalization statistics via a moving average update, de-noises the embedding space, and facilitates TD to reconstruct cleaner images. Our method enables the robot to navigate using these reconstructed images as surrogate observations instead of the actual observations. TD is trained via self-supervised learning on “uncorrupted” images to learn a representation capable of reconstructing input images. Interestingly, TD can reconstruct cleaner images given corrupted images during the test phase.
@@
In summary, we develop a benchmark consisting of 13 types of visual corruptions to evaluate how navigation performance is degraded by each corruption type, as detailed in Fig. 3. These challenges highlight the limitations of non-adaptive models in maintaining effective navigation under visual corruptions. Here, “non-adaptive” refers to models that do not dynamically adjust their processing or decision-making mechanisms in response to changes in visual input quality. Then, we propose a TTA method for point-goal navigation, named TTA-Nav. It is a “plug-and-play” method as it can be seamlessly integrated into existing navigation models without re-training. TTA-Nav leverages a top-down decoder to reconstruct cleaner versions of corrupted images. Our method improves navigation performance without explicitly fine-tuning the decoder on corrupted inputs. It outperforms three state-of-the-art TTA methods: DUA [11], TENT [12], and SHOT-IM [13]. This is evidenced across our visual corruption benchmarks, where TTA-Nav consistently shows improved performance.
@@ III-A Point-Goal Navigation Problem Formulation
In point-goal navigation, a robot is tasked to move to the given relative direction: $(\Delta x_{\text{target}},\;\Delta y_{\text{target}})=(x_{\text{target}}-x_{\text{start}},y_{\text{target}}-y_{\text{start}})$ in a cluttered indoor scene. The robot takes both RGB frames and perfect odometry as input. For instance, given the goal of moving to $(\Delta x_{\text{target}}=0.5m,\Delta y_{\text{target}}=-2m)$, the robot has to move to the target position $(x_{\text{start}}+0.5,y_{\text{start}}-2)$, where $(x_{\text{start}},y_{\text{start}})$ is the initial position. The short path will lead to the high reward given to the robot. In our setting, the robot is equipped with a navigation model trained with clean RGB observations and the perfect odometry $(\Delta x_{t},\Delta y_{t})$. Where $\Delta x_{t}=x_{t}-x_{t-1}$, $\Delta y_{t}=y_{t}-y_{t-1}$, and ($x_{t}$, $y_{t}$) is the position of the robot at time $t$. The robot is required to navigate in unseen scenes with visual corruptions. We use the Gibson scene dataset [39] by following the standard train and validation split as in [23] which yields 72 training and 14 validation scenes. The agent is trained with training scenes and evaluated with 14 unseen validation scenes (994 episodes). We use Habitat platform [20, 40, 41] as an environment. Specifically, the agent has access to an RGB observation ($256\times 256$) and GPS & Compass $(x_{\text{target}}-x_{t},y_{\text{target}}-y_{t})$, and can take 4 actions: move forward (0.25m), turn left (10°), turn right (10°) and stop. GPS & Compass is simply derived from the odometry, given a relative target position at start. Once the agent takes stop action, an episode will end. During the evaluation, we allow the agent to take up to 500 environment steps in an episode, i.e., timed-out steps.
@@ III-C Algorithm
Training: During training of TTA-Nav, we freeze all VE parameters and train TD with mean squared error (MSE) reconstruction loss with samples from 72 training scenes of Gibson [39] without visual corruption. TD is trained with a very small subset of navigation data: 112k out of 2.5 billion frames. We follow the same optimization hyperparameters as in [47]. Specifically, we use ADAM optimizer [48] with the learning rate of $2\times 10^{5}$, and the momentum of $0.9999$. See Algorithm 1 for more details. The code will be made publicly available upon publication.
Adaptation: TTA-Nav allows batch statistics in the BatchNorm layers of the VE to be updated during the test phase. Specifically, each batch normalization layer normalizes an incoming sample $x$ by the estimated mean and variance of the training distribution as in (1),
@@
where $\hat{\mu}_{k}$ is the estimated mean and $\hat{\sigma}^{2}_{k}$ is the estimated variance of the training distribution, $\mu_{k}$ and $\sigma^{2}_{k}$ are the sample mean and variance of the batch data, and $\rho$ is a hyperparameter called momentum, which determines the contribution of the most recent sample estimate ($\mu_{k},\sigma^{2}_{k}$) to the calculation of the average. Here we update $\mu_{k}$ and $\sigma_{k}$ at every environment step, so the batch size is 1. See Algorithm 2 for more detail. In subsection IV-D, we will show that this simple adaptation can restore clean images from visual corruption.
@@ IV Experiments
We will describe the experiment setup which includes visual corruptions, baseline methods, and evaluation metrics. Subsequently, we will show experimental results in the following sequence. Firstly, we examine the visual content reconstructed by TD to illustrate why TTA-Nav is effective for navigation under visual corruptions. Secondly, we compare TTA-Nav with the baseline methods on clean and corrupted observations. Lastly, we conduct an ablation study to see what components lead to improvement in performance. All models in all experiments are first trained on clean observations and then adapt to changing environments with different corruptions during the testing phase.
@@ IV-B Baselines
  • Pretrained-Nav is our navigation model that is fine-tuned from DD-PPO as explaineded in section III.
@@ IV-D Results
Normalization implicitly does Image Restoration: We compare reconstructions from test input images by TD with and without updating normalization parameters. Qualitatively, we can see in Fig. 4 that the reconstruction from adapted TD is brighter when input is dimmed and clearer when input is blurred or noisy. Color is changed to normal when the input image is dyed with yellow. Some water drops are removed by adapted TD. Moreover, if the input is clean, we can barely see the difference between the reconstructions of adapted and non-adapted TD.
@@ V Conclusion
We present Test-time Adaptation (TTA) for point-goal navigation through our method TTA-Nav. Its key ideas are Top-down Decoder (TD) and adaptive normalization in a pre-trained navigation model. TD is integrated seamlessly with the pre-trained navigation model, offering robust adaptability to visually corrupted conditions without explicitly fine-tuning. TTA-Nav enhances navigation performance under visual corruptions without gradient-based optimization during the test time. Experimental results showcase TTA-Nav’s superior performance compared to SOTA TTA methods, including DUA [11], TENT [12], and SHOT-IM [13]. The results underscore its potential for widespread applications in robot navigation and other robotic tasks.

### 14 Mar 2024 v2

@@ Abstract
Robot navigation under visual corruptions presents a formidable challenge. To address this, we propose a Test-time Adaptation (TTA) method, named as TTA-Nav, for point-goal navigation under visual corruptions. Our “plug-and-play” method incorporates a top-down decoder to a pre-trained navigation model. Firstly, the pre-trained navigation model receives a corrupted image input and extracts features. Secondly, the top-down decoder produces the reconstruction given the high-level features extracted by the pre-trained model. Then, the decoder feeds the reconstruction of a corrupted image back to the pre-trained model, which outputs the final action. Despite being trained solely on clean images, the top-down decoder reconstructs cleaner images from corrupted ones without gradient-based adaptation. The pre-trained navigation model with our top-down decoder significantly enhances navigation performance across almost all visual corruptions in our benchmarks. Our method improves the success rate of point-goal navigation from the state-of-the-art result of 46% to 94% on the most severe corruption, showing the potential for broader application in robotic visual navigation. Project page: https://sites.google.com/view/tta-nav
@@ I Introduction
Robot navigation faces substantial challenges when confronted with visual corruptions [1, 2, 3, 4, 5], as illustrated in Fig. 1: The robot is equipped with a state-of-the-art (SOTA) navigation model trained on 2.5 billion clean frames with deep reinforcement learning (DRL). Despite being trained on a great amount of data, the robot still fails to navigate to the target position under dimmed lighting. Unforeseeable visual corruptions can occur during the test time and these corruptions are harmful to navigation performance.
@@
This challenge raises questions about the adaptability of robots: How can a robot dynamically adjust in real-time to visual corruptions that are unforeseeable in the training phase? And how can it maintain effective navigation when encountering such corruption despite being trained solely on clean observations? The previous work [5] considers the navigation with a map, which can be deteriorated if the camera images are corrupted, and proposes a map refiner to generate a more precise map. However, their method requires the robot to collect data and gradient-based fine-tuning in the test time. That means the robot has a distinct fine-tuning and test phase. This is impractical in some applications that require the robot to immediately adapt to corruption, such as self-driving cars.
@@
Figure 1: Point-Goal Navigation Under Visual Corruptions: A robot is tasked to move from start to goal position. Top: The robot is equipped with the state-of-the-art navigation method [6]. It fails to navigate from the bedroom to another room when facing the dimmed light condition. Bottom: The robot is equipped with our method. It receives the reconstructed image as the “surrogate observation” instead of the real observation from the scene.
@@
Figure 2: An overview of our proposed method (TTA-Nav): We present the pretrained navigation model (Pretrained-Nav), which contains Visual Encoder (VE) and policy network, in the purple frame. TTA-Nav is a plug-and-play method. It has Top-down Decoder (TD) in the orange frame. TD receives the output of the late layer of VE and projects the reconstructed image back to the VE’s input layer. Here, VE is SE-ResNeXt-50 [7] and TD receives block3’s output. Left: During training, VE remains fixed, while TD is trained to predict VE’s input by minimizing Mean Squared Error (MSE) loss. Training occurs offline, utilizing samples from Replay, which contains the visual experiences of a robot navigating 72 Gibson training scenes. Right: During testing, TD is frozen and the BatchNorm layers of VE can adjust their normalization statistics $\hat{\mu}_{k}$ and $\hat{\sigma}_{k}$. See section III for more details. Finally, the reconstructed image is input to VE for navigation. VE does feedforward computation again and the policy outputs actions.
@@
To this end, we propose a Test-time Adaptation (TTA) method for robot navigation under visual corruptions. Our method is inspired by the top-down processing observed in human cognition [8]. This processing is deemed as a fundamental mechanism for solving challenging visual tasks [9, 10]. Specifically, consider a Deep Neural Network (DNN) model as an architecture with low-level visual processing at earlier layers and high-level visual processing at later layers, our method relies on top-down modulation from a decoder, which receives the high-level features of an image extracted by the pre-trained visual perception model, reconstructs the clean observation and feeds reconstructed inputs back to the visual perception model. The other key mechanism of our method is adaptive normalization (AN) in the pre-trained visual perception model. During the test phase, AN adjusts normalization statistics via a moving average update, de-noises image features, and enables the decoder to reconstruct cleaner images. Our method enables the robot to navigate using these reconstructed images as surrogate observations instead of the actual observations. The decoder is trained via self-supervised learning on “uncorrupted” images to learn a representation capable of reconstructing input images. Interestingly, the decoder can reconstruct cleaner images given corrupted images during the test phase.
@@
In summary, we develop a benchmark consisting of 13 types of visual corruptions to evaluate how navigation performance degrades under each corruption type, as detailed in Fig. 3. These challenges highlight the limitations of non-adaptive models in maintaining effective navigation under visual corruptions. Here, “non-adaptive” refers to models that do not dynamically adjust their processing or decision-making mechanisms in response to changes in visual input quality. Then, we propose a TTA method for point-goal navigation, named TTA-Nav, which can be seamlessly integrated, i.e., “plug-and-play”, into existing navigation models without re-training. Particularly, TTA-Nav leverages adaptive normalization in the pre-trained visual perception model to denoise image features, and uses a top-down decoder to reconstruct cleaner versions of corrupted images from the high-level features. Our method improves navigation performance without explicitly fine-tuning the decoder on corrupted inputs. It outperforms three state-of-the-art TTA methods: DUA [11], TENT [12], and SHOT-IM [13]. This is evidenced across our visual corruption benchmarks, where TTA-Nav consistently shows improved performance.
@@ III-A Point-Goal Navigation Problem Formulation
In point-goal navigation, a robot is tasked to move to the given relative direction: $(\Delta x_{\text{target}},\;\Delta y_{\text{target}})=(x_{\text{target}}-x_{\text{start}},y_{\text{target}}-y_{\text{start}})$ in a cluttered indoor scene. For instance, given the goal of moving to $(\Delta x_{\text{target}}=0.5m,\Delta y_{\text{target}}=-2m)$, the robot has to move to the target position $(x_{\text{start}}+0.5,y_{\text{start}}-2)$, where $(x_{\text{start}},y_{\text{start}})$ is the initial position. The short path will lead to the high reward given to the robot. In our setting, the robot is equipped with a navigation model trained with clean RGB observations and GPS & Compass $(x_{\text{target}}-x_{t},y_{\text{target}}-y_{t})$. The GPS & Compass is derived from the error-free odometry $(\Delta x_{t},\Delta y_{t})$. Where $\Delta x_{t}=x_{t}-x_{t-1}$, $\Delta y_{t}=y_{t}-y_{t-1}$, and ($x_{t}$, $y_{t}$) is the position of the robot at time $t$. The robot can take 4 actions: move forward (0.25m), turn left (10°), turn right (10°) and stop. Once the robot takes stop action, an episode will end. During the evaluation, we allow the robot to take up to 500 environment steps in an episode, i.e., timed-out steps. We use the Gibson scene dataset [39] by following the standard train and validation split as in [23] which yields 72 training and 14 validation scenes. The robot is trained with training scenes and evaluated with 14 unseen validation scenes (994 episodes). We use Habitat platform [20, 40, 41] as an environment.
@@ III-C Algorithm
Training: During the training of TTA-Nav, we freeze all VE parameters and train TD with mean squared error (MSE) reconstruction loss with samples from 72 training scenes of Gibson [39] without visual corruption. TD is trained with a very small subset of navigation data: 112k out of 2.5 billion frames. We follow the same optimization hyperparameters as in [47]. Specifically, we use ADAM optimizer [48] with the learning rate of $2\times 10^{5}$, and the momentum of $0.9999$. See Algorithm 1 for more details.
Adaptation: TTA-Nav allows batch statistics in the BatchNorm layers of the VE to be updated during the test time. We call this test-time update adaptive normalization. Specifically, each batch normalization layer normalizes an incoming sample $x$ by the estimated mean and variance as in (1),
@@
where $\hat{\mu}_{k}$ is the estimated mean and $\hat{\sigma}^{2}_{k}$ is the estimated variance, $\mu_{k}$ and $\sigma^{2}_{k}$ are the sample mean and variance of the batch data, and $\rho$ is a hyperparameter called momentum, which determines the contribution of the most recent sample estimate ($\mu_{k},\sigma^{2}_{k}$) to the calculation of the average. Here we update $\mu_{k}$ and $\sigma_{k}$ at every environment step, so the batch size is 1. See Algorithm 2 for more detail. In subsection IV-D, we will show that this simple adaptation can restore clean images from visual corruption.
@@ IV Experiments
We will describe the experimental setup which includes visual corruptions, baseline methods, and evaluation metrics. Subsequently, we will show experimental results in the following sequence. Firstly, we examine the visual content reconstructed by TD to illustrate why TTA-Nav is effective for navigation under visual corruptions. Secondly, we compare TTA-Nav with the baseline methods on clean and corrupted observations. Lastly, we conduct an ablation study to see what components lead to improvement in performance. All models in all experiments are first trained on clean observations and then adapt to changing environments with different corruptions during the testing phase.
@@ IV-B Baselines
  • Pretrained-Nav is our navigation model that is fine-tuned from DD-PPO as explained in section III.
@@ IV-D Results
Adaptive Normalization implicitly does Image Restoration: We compare reconstructions from test input images by TD with and without updating normalization parameters. Qualitatively, we can see in Fig. 4 that the reconstruction from adapted TD is brighter when input is dimmed and clearer when input is blurred or noisy. Color is changed to normal when the input image is dyed with yellow. Some water drops are removed by adapted TD. Moreover, if the input is clean, we can barely see the difference between the reconstructions of adapted and non-adapted TD.
@@ V Conclusion
We present Test-time Adaptation (TTA) for point-goal navigation through our method TTA-Nav. Its key ideas are Top-down Decoder (TD) and adaptive normalization in the pre-trained navigation model. TD is integrated seamlessly with the pre-trained navigation model, offering robust adaptability to visually corrupted conditions without explicitly fine-tuning. TTA-Nav enhances navigation performance under visual corruptions without gradient-based optimization during the test time. Experimental results showcase TTA-Nav’s superior performance compared to SOTA TTA methods, including DUA [11], TENT [12], and SHOT-IM [13]. The results underscore its potential for widespread applications in robot navigation and other robotic tasks.